OpenStudy (stuck-help):

METAL square root r^2+49=49

OpenStudy (jinx.exe):

Simplifying r2 + 49 = 49 Reorder the terms: 49 + r2 = 49 Add '-49' to each side of the equation. 49 + -49 + r2 = 49 + -49 Combine like terms: 49 + -49 = 0 0 + r2 = 49 + -49 r2 = 49 + -49 Combine like terms: 49 + -49 = 0 r2 = 0 Solving r2 = 0 Solving for variable 'r'. Move all terms containing r to the left, all other terms to the right. Simplifying r2 = 0 Take the square root of each side: r = {0}